B2 Consider the following information on a portfolio of three stocks: State of Probabil f Stock A Stock B Stock C Rate of Return Rate of Return Rate of Return 21 15 -22 Economy State of Ec Boom Normal Bust 15 80 05 .05 08 18 .07 The portfolio is invested 35 percent in each Stock A and Stock B and 30 percent in Stock C If the expected T-bill rate is 3.90 percent, what is the expected risk premium on the portfolio? (10 marks) i) Explain the relationships among the reward-to-risk ratio, risk-free rate of return, market rate of return, market risk premium, beta, and the security market line. (10 marks) (20 marks
Answer #1

ANSWER (I)

E(r)Boom= (0.35 *0.05) + (0.35 *0.21) + (0.30 *0.18) = 0.1450

E(r)Normal= (0.35 *0.08) + (0.35 *0.15) + (0.30 *0.07) = 0.1015

E(r)Bust= (0.35 *0.12) + (0.35 *-0.22) + (0.30 *-0.02) = -0.0410

And the expected return of the portfolio is:

E(r)P= (0.15 *0.1450) + (0.80 *0.1015) (0.05 *-0.0410) = 0.1009

The risk premium is the return of a risky asset, minus the risk-free rate. T-bills are often used as the risk-free rate,so

ERPP= 0.1009 - 0.039 = 6.19 %

----------------------------------------------------------------------

ANSWER (ii)

The reward-to-risk ratio is the risk premium of market and is the slope of the security market line whereas risk-free rate is the vertical intercept of the line.

The market rate of return is present rate prevailing in the market that is interest for any amount accepted by the borrowers this is assigned a beta of1.0. All securities that are promptly priced based on the capital asset pricing model, will plot on the security market line at a location determined by the security's beta